OpenStudy (anonymous):

factoring perfect square trinomials and differences of squares...factor completely, remember to look 1st for a common factor. check by multiplying. if it is a prime state it. 9x^2+30x+25

OpenStudy (anonymous):

9x^2+30x+25 = 9x^2 + 15x + 15x + 25 = 3x(3x + 5) +5(3x + 5) = (3x + 5)(3x + 5)

OpenStudy (anonymous):

how do i check my answer by multiplying? Is it a prime? How do you know if it is a prime or not?

OpenStudy (anonymous):

We can also solve it using the identity a^2 + 2ab + b^2 = (a + b)^2 9x^2+30x+25 = (3x)^2 + 2(3x)(5)+ (5)^2 using the identity a^2 + 2ab + b^2 = (a + b)^2 = (3x + 5)^2 so 9x^2+30x+25 = (3x + 5)^2 or 9x^2+30x+25 = (3x = 5)(3x + 5)

OpenStudy (anonymous):

Prime factor is a factor from which nothing more can be taken out as common suppose we get something like (3x - 9) It is not prime as we can take out 3 as common factor i.e. 3(x - 3) In above case we have got both factors as 3x + 5 there is nothing common between 3x and 5......... so they r prime factors

OpenStudy (anonymous):

so to check my answer i would multiply the final equation, correct? 3x*3x=9x^2 3x*5=15x 5*3x=15x 5*5=25 combine like terms- 15x/15x=30x 9x^2+30x+25

OpenStudy (anonymous):

Yes that is correct.....

OpenStudy (anonymous):

what is the greates common factor? Is it 3?